# Accretion onto supermassive and intermediate mass black holes in cosmological simulations

### v1

@@ Acknowledgements.
RW acknowledges funding of a Leibniz Junior Research Group (project number J131/2022). This work was supported by the Simons Collaboration “Learning the Universe”. R.W. acknowledges support from the NSF via XSEDE allocation PHY210011 in the early phases of this project. L.B. acknowledges support from NASA award #80NSSC22K0808 and NSF award AST-2307171. G.L.B. acknowledges support from the NSF (AST-2108470 and AST-2307419, ACCESS), a NASA TCAN award, and the Simons Foundation through the Learning the Universe Collaboration.References
@@
- Guo et al. (2024) Guo, M., Stone, J. M., Quataert, E., & Kim, C.-G. 2024, arXiv e-prints, arXiv:2405.11711
@@
- Li et al. (2024) Li, J., Silverman, J. D., Shen, Y., et al. 2024, arXiv e-prints, arXiv:2403.00074
@@
- Maiolino et al. (2023) Maiolino, R., Scholtz, J., Curtis-Lake, E., et al. 2023, arXiv e-prints, arXiv:2308.01230
@@
- Martín-Navarro et al. (2018) Martín-Navarro, I., Brodie, J. P., Romanowsky, A. J., Ruiz-Lara, T., & van de Ven, G. 2018, Nature, 553, 307
@@ Appendix A Impact on the galaxy formation model
As MBHs have a substantial impact on the evolution of their host galaxies (Fabian 2012), it is reasonable to assume that changing the gas accretion formula has an impact on the galaxy population. We explore this using the calibration plots in IllustrisTNG (Pillepich et al. 2018), and show the median trends of stellar mass fraction vs. halo mass, gas fraction vs. halo mass, star formation rate density vs. redshift, stellar half mass radius vs. stellar mass, stellar mass function and B-V colour vs. stellar mass for the different models in Figure 10. While there are some notable differences in all of these plots, we argue that they originate from the same single cause: the precise mass scale where efficient AGN feedback sets in.
@@
where denotes a characteristic black hole mass scale (Weinberger et al. 2017). This leads to a clear transition with less massive black holes predominantly powering inefficient thermal feedback and more massive black holes predominantly powering efficient kinetic feedback (Weinberger et al. 2018). Consequently, the stellar and halo mass scale at which AGN feedback becomes efficient depends strongly on the mass of the black holes in them, and most of the differences in the resulting scaling relation can be tied back to the changed black hole mass - stellar mass and black hole mass - halo mass relations.
Considering the black hole mass - stellar mass relation of Figure 5 in more detail, the black holes reach the mass of at higher halo and stellar mass in the ff model, and at lower masses in the mod ff model (compared to the tng model). This leads to massive galaxies in the ff model remaining star forming up to higher stellar and halo masses, leading to higher stellar mass fraction in the high-mass end (top left panel of Figure 10). The lack of feedback implies the retention of higher gas fractions at halo masses around , and the additional star formation in these systems shows up as slightly higher star formation rate densities at low redshift. The later shutoff of star formation in massive galaxies leads to fewer dry mergers that could increase stellar half mass radii in high mass galaxies and consequently more compact galaxies in this regime, as well as a transition from red to blue colours happening at higher masses. For the galaxies in the mod ff model, the efficient AGN feedback sets in at lower halo masses, as clearly visible in the drop in gas fractions above . This leads to a slight drop in star formation rate density toward redshift 0, as well as a transition to red, quiescent galaxies at a lower stellar mass. It leads to a slight reduction in stellar mass fraction and increase in stellar half mass radii, likely due to the increased dry-merger rate dominating the stellar mass evolution in this regime over in-situ star formation. Thus, for both inflow rate variations, simply adjusting the transition mass would likely compensate for the effects on the galaxy population. We note there are some other minor changes in the low-mass population, however, we refrain from over-interpreting these due to the moderate resolution simulations used in this work. Finally, the unresolved accretion disk models have a negligible effect on the galaxy population, similar to their role in the global black hole mass density buildup over cosmic time.
@@ Appendix C Numerical convergence
To study the numerical convergence of the new methods, we performed each simulation at resolution levels (see Table 1). We show black hole accretion rate density and black hole mass density relative to the highest resolution run in Figure 12 for the L50, SMBH runs and Figure 13 for the L12.5, IMBH runs. While the accretion rates mostly increase with higher resolution, the resolution effects are relatively moderate, except for the lowest resolution L50 runs.

### v2

@@ Acknowledgements.
We thank the anonymous referee for the suggestions that helped to improve the manuscript. RW acknowledges funding of a Leibniz Junior Research Group (project number J131/2022). This work was supported by the Simons Collaboration “Learning the Universe”. R.W. acknowledges support from the NSF via XSEDE allocation PHY210011 in the early phases of this project. L.B. acknowledges support from NASA award #80NSSC22K0808 and NSF award AST-2307171. G.L.B. acknowledges support from the NSF (AST-2108470 and AST-2307419, ACCESS), a NASA TCAN award, and the Simons Foundation through the Learning the Universe Collaboration.References
@@
- Cochrane et al. (2023) Cochrane, R. K., Anglés-Alcázar, D., Mercedes-Feliz, J., et al. 2023, MNRAS, 523, 2409
@@
- Farcy et al. (2025) Farcy, M., Hirschmann, M., Somerville, R. S., et al. 2025, arXiv e-prints, arXiv:2504.08041
@@
- Gordon et al. (2024) Gordon, S. T., Smith, B. D., Khochfar, S., & Regan, J. A. 2024, MNRAS, 529, 604
@@
- Guo et al. (2024) Guo, M., Stone, J. M., Quataert, E., & Kim, C.-G. 2024, ApJ, 973, 141
@@
- Li et al. (2025) Li, J., Silverman, J. D., Shen, Y., et al. 2025, ApJ, 981, 19
@@
- Maiolino et al. (2024) Maiolino, R., Scholtz, J., Curtis-Lake, E., et al. 2024, A&A, 691, A145
@@
- Mart´ın-Navarro et al. (2018) Martín-Navarro, I., Brodie, J. P., Romanowsky, A. J., Ruiz-Lara, T., & van de Ven, G. 2018, Nature, 553, 307
@@
- Ostriker et al. (2010) Ostriker, J. P., Choi, E., Ciotti, L., Novak, G. S., & Proga, D. 2010, ApJ, 722, 642
@@ Appendix A Impact on the galaxy formation model
As MBHs have a substantial impact on the evolution of their host galaxies (Fabian 2012), it is reasonable to assume that changing the gas accretion formula has an impact on the galaxy population. We explore this using the calibration plots in IllustrisTNG (Pillepich et al. 2018), and show the median trends of stellar mass fraction versus halo mass, gas fraction versus halo mass, star formation rate density versus redshift, stellar half mass radius versus stellar mass, stellar mass function and B-V colour versus stellar mass for the different models in Figure 10. While there are some notable differences in all of these plots, we argue that they originate from the same single cause: the precise mass scale where efficient AGN feedback sets in.
@@
where denotes a characteristic BH mass scale (Weinberger et al. 2017). This leads to a clear transition with less massive BHs predominantly powering inefficient thermal feedback and more massive BHs predominantly powering efficient kinetic feedback (Weinberger et al. 2018). Consequently, the stellar and halo mass scale at which AGN feedback becomes efficient depends strongly on the mass of the BHs in them, and most of the differences in the resulting scaling relation can be tied back to the changed BH mass-stellar mass and BH mass-halo mass relations.
Considering the BH mass-stellar mass relation of Figure 5 in more detail, the BHs reach the mass of at higher halo and stellar mass in the ff model, and at lower masses in the mod ff model (compared to the tng model). This leads to massive galaxies in the ff model remaining star forming up to higher stellar and halo masses, leading to higher stellar mass fraction in the high-mass end (top left panel of Figure 10). The lack of feedback implies the retention of higher gas fractions at halo masses around , and the additional star formation in these systems shows up as slightly higher star formation rate densities at low redshift. The later shutoff of star formation in massive galaxies leads to fewer dry mergers that could increase stellar half mass radii in high mass galaxies and consequently more compact galaxies in this regime, as well as a transition from red to blue colours happening at higher masses. For the galaxies in the mod ff model, the efficient AGN feedback sets in at lower halo masses, as clearly visible in the drop in gas fractions above . This leads to a slight drop in star formation rate density toward redshift 0, as well as a transition to red, quiescent galaxies at a lower stellar mass. It leads to a slight reduction in stellar mass fraction and increase in stellar half mass radii, likely due to the increased dry-merger rate dominating the stellar mass evolution in this regime over in-situ star formation. Thus, for both inflow rate variations, simply adjusting the transition mass would likely compensate for the effects on the galaxy population. We note there are some other minor changes in the low-mass population, however, we refrain from over-interpreting these due to the moderate resolution simulations used in this work. Finally, the unresolved accretion disk models have a negligible effect on the galaxy population, similar to their role in the global BHMD buildup over cosmic time.
@@ Appendix C Numerical convergence
To study the numerical convergence of the new methods, we performed each simulation at resolution levels (see Table 1). We show the BHARD and BHMD relative to the highest resolution run in Figure 12 for the L50, SMBH runs and for the L12.5, IMBH runs in the left and right panels, respectively. While the accretion rates mostly increase with higher resolution, the resolution effects are relatively moderate, except for the lowest resolution L50 runs.
Appendix D Environment of IMBHs
To illustrate the environment of IMBHs at high redshift, Figure 13 shows the gas column density in halo 0 at redshift 5 and its main progenitor at redshift 6, 8 and 10. The dotted, dashed and solid circle show , and the accretion radius, respectively. While the accretion radius at approaches the galaxy radius, the gas structure at these early times is clearly limited by the mass resolution and modelling limitations.